OpenStudy (anonymous):

Discrete. I am trying to get better at proofs, especially the notation. A product of any two irrational numbers is irrational.

OpenStudy (anonymous):

So I would like to let statement r1=(Let m = sqrt(2) and Let n = sqrt(2)

OpenStudy (anonymous):

Therefore r1 implies r2 with contradicts our original statement. So The initial statement " A product of any two irrationals," is False.

OpenStudy (anonymous):

Going for proof by contradiction...

jimthompson5910 (jim_thompson5910):

that's one counter example another is let m = sqrt(3) and n = sqrt(12) m & n are definitely irrational, but m*n = sqrt(3)*sqrt(12) = sqrt(3*12) = sqrt(36) = 6 which shows that m*n is rational

OpenStudy (anonymous):

So should I try to prove by cases and show multiple? Or is there a way I can just do one example for proof by contradiction?

jimthompson5910 (jim_thompson5910):

once you show one counterexample, it disproves the whole statement

jimthompson5910 (jim_thompson5910):

because the claim is for ALL irrational numbers (or any two irrationals)

jimthompson5910 (jim_thompson5910):

find one hole and the whole thing falls apart

OpenStudy (anonymous):

Thank you Jim!